# DECIPHERING THE DEFINITION OF ADVERSARIAL ROBUSTNESS FOR POST-HOC OOD DETECTORS

### v3

@@ Abstract
Detecting out-of-distribution (OOD) inputs is critical for safely deploying deep learning models in real-world scenarios. In recent years, many OOD detectors have been developed, and even the benchmarking has been standardized, i.e. OpenOOD. The number of post-hoc detectors is growing fast and showing an option to protect a pre-trained classifier against natural distribution shifts, claiming to be ready for real-world scenarios. However, its efficacy in handling adversarial examples has been neglected in the majority of studies. This paper investigates the adversarial robustness of 16 post-hoc detectors on several evasion attacks and discusses a roadmap toward adversarial defense in OOD detectors.
@@ 1 INTRODUCTION
Adversarial robustness in the context of out-of-distribution (OOD) detection refers to the ability of a detector to correctly identify OOD samples even when they have been adversarial perturbed to evade a deep neural network (DNN). Evasion attacks, e.g. PGD [1], which are designed to fool deep learning classifiers, are difficult to spot as an outlier for OOD detectors. To prevent errors in real-world applications, it is crucial to detect OOD cases, not only for natural distribution shifts [2] but also for adversarial examples [3] without degrading the generalizability of the underlying pre-trained classifier [4].
Current standardized benchmarks such as OpenOOD [5] and RoboDepth [6] merely focus on natural distribution shifts and corruptions [7]. Especially OpenOOD aims to make a fair comparison across methods initially developed for anomaly detection [8], model uncertainty [9], open-set recognition (OSR) [10], and OOD detection [11] but neglecting adversarial examples.
OpenOOD benchmark suite [4] evaluates methods on semantic shift (e.g., samples that are semantically different from the training data, representing truly novel or unseen concepts.) [12] and covariate shift (e.g. samples that come from a different distribution than the training data, but still belong to the same semantic categories).
@@
Current OOD detection methods, as some listed in Table 1, achieve outstanding results on prominent OOD benchmarks, such as the OpenImage-O [13], ImageNet-O [14], Texture [15], and iNaturalist [16, 17]. OOD detection is a very quickly growing field by the number of methods added to OpenOOD. More specific post-hoc methods with their plug-and-play capabilities on pre-trained classifiers are more flexible and scalable compared to methods that require full retraining on new OOD data [18, 19]. Simple post-hoc methods like KNN [20] are highlighted maintaining good performance on toy-datasets (e.g. MNIST [21], CIFAR-10 or CIAR-100 [22]), and also show outstanding performance on the more realistic dataset like ImageNet [23] according to [18]. These experiments do neglect the adversarial robustness of “state-of-the-art” (sota) detectors and the real-world capabilities [24] are questionable as past studies had shown [3, 25, 26, 27]. OOD detectors should be able to reliably detect inputs that come from a different distribution than the training data, even if the shift is subtle or complex. Adversarial examples remain challenging because they share the same semantics as the training data, but aim to modify the classifier’s output.
@@
  • We revise the definition of adversarial OOD regarding post-hoc OOD methods to finally have a common understanding of robust adversarial OOD detection.
  • We examine 16 post-hoc OOD detectors by delving into their current ability to detect adversarial examples — an aspect that has been disregarded.
  • We expand the OpenOOD framework with evasive attacks and provide adversarial OOD datasets: github.com/adverML/AdvOpenOOD.
@@ 2 RELATED WORK
This section discusses existing white-box evasion attacks, the benefits of post-hoc detectors, existing methods for detecting adversarial samples, and finally Grad-CAM, a tool for understanding the neural networks’ attention shift initiated by adversarial examples.
@@ 2.1 Evasion Attacks Crafting Inliers
The objective of evasion attacks is to generate adversarial examples that will result in the misclassification of inputs at deep learning models [28]. It is possible to distinguish between two types of attack: black-box attacks [29], where the classifier is queried, and white-box attacks [30], where the network is under the attacker’s complete control. A white-box threat model is strictly stronger. They try to find the smallest possible perturbation, often imperceptible to humans, to manipulate the model’s decision boundaries.
More formally, for an input ${\bm{x}}$ with the ground-truth label $y$, an adversarial example ${\bm{x}}^{\prime}$ is crafted by adding small noise $\bm{\delta}$ to ${\bm{x}}$ such that the predictor model loss ${\bm{J}}({\bm{x}}^{\prime},y)$ is maximized. The $L^{p}$ norm of the adversarial noise should be less than a specified value $\epsilon$, i.e., $\left\|{\bm{x}}-{\bm{x}}^{\prime}\right\|\leq\epsilon$, e.g., $\epsilon=8/255$ [31], to ensure that the image does not change semantically. The attack method Fast Gradient Sign Method (FGSM) by [32] maximizes the loss function in a single step by taking a step towards the sign of the gradient of ${\bm{J}}({\bm{x}},y)$ w.r.t. to ${\bm{x}}$: ${\bm{x}}^{\prime}={\bm{x}}+\epsilon\text{sign}(\nabla_{{\bm{x}}}{\bm{J}}({\bm{x}}^{\prime},y))$, where the noise meets the $L^{\infty}$ norm bound $\epsilon$. Furthermore, this approach can be applied iteratively, as shown by [33], using a reduced step size $\alpha$: $vx^{\prime}_{0}={\bm{x}},\;{\bm{x}}^{\prime}_{t+1}={\bm{x}}^{\prime}_{t}+\alpha\epsilon\text{sign}(\nabla_{{\bm{x}}}{\bm{J}}({\bm{x}}^{\prime}_{t},y))$, where in each step, the perturbation should be constrained within the $L^{\infty}$ ball of radius $\epsilon$. This constraint is a characteristic of the Projected Gradient Descent (PGD) attack proposed by [1], which is commonly considered a standard attack for the evaluation of model robustness [34]. Masked PGD (mPGD) [35] is a variant of the PGD attack that restricts perturbations to a specific region within an image. The PGD attack is extended towards a patch: ${\bm{x}}_{t+1}^{\prime}=\text{Clip}_{{\bm{x}}^{\prime}}\left({\bm{x}}_{t}^{\prime}+\alpha\cdot\text{sign}(\nabla_{{\bm{x}}}{\bm{J}}({\bm{x}}_{t}^{\prime},y,\bm{\theta})[patch]),\epsilon\right).$ In this context, the term “patch” refers to the region specified as $[x:x+h,y:y+w]$, where $[x,y,h,w]$ are the provided patch’s coordinates and dimensions.
@@ 2.2 Advantages of Post-Hoc OOD Detectors
Moreover, post-hoc OOD methods can be augmented with other techniques, such as those employed in OSR [40] or uncertainty estimation [41]. Combining different techniques means also the post-hoc methods become more complex and thus the attack surface might become larger, i.e. attacks against uncertainty estimation [9] differ to the evasion attacks. It is not necessarily the case that a post-hoc method must be combined with other techniques. This demonstrates SAFE [42].
@@ 2.3 OOD Adversarial Detection
Ensuring the protection of deployed DL models is the aim of OOD detectors, but the task of providing comprehensive defense [43] against unknown threats is challenging. Every defense mechanism can be circumvented at some point [44]. Many OOD detectors can be easily evaded by slightly perturbing benign OOD inputs, creating OOD adversarial examples that reveal a severe limitation of current open-world learning frameworks [45, 46]. Even adversarial training-based defense methods, effective against ID adversarial attacks, struggle against OOD adversarial examples [46]. In the past years, various defensive techniques [47] and combinations of them have surfaced to combat threats such as adversarial training [1, 48, 49], gradient masking for obfuscation [50], input transformations such as input purification [51, 52]. However, attackers consistently adapt their adversarial attacks to the specific defense mechanisms [53]. According to [54], optimization-based defenses could be a promising future, because they can adapt during test-time towards the input.
@@
2.4 CAM-based Explanations and Consistent Shift by Adversarial Examples
Explainable AI (XAI) [57] can be used as a defense mechanism [58, 59, 60] to spot adversarial examples [61]. The Grad-CAM heatmap provides a clear indication, for a given image, of the regions that the Convolutional Neural Network (CNN) focuses on to make a decision. Class Activation Maps (CAMs) can be considered saliency maps specific to the input [62]. These maps emerge from the accumulated and up-scaled activations at a particular convolutional layer, typically the penultimate layer. The classification is estimated through a linear combination of the activation of units in the final layer(s) of the feature selection network:
$f_{\theta}(.)\approx\sum_{i}\sum_{k}w_{k}a_{ki},$ (1)
where $a_{ki}$ represents the activation of the k-th channel of unit $i$, and $w_{k}$ denotes the learned weights. The relevance values are subsequently defined as $r_{i}=\sum_{k}w_{k}a_{ki}$. The determination of these weights depends on the specific CAM variant employed [63, 64, 65]. Grad-CAM weights the activations using gradients:
$w_{K}:=\frac{\partial f_{\theta}(.)}{\partial a_{ki}}.$ (2)
This weighting directly links to more fundamental explanations that merely estimate the influence of the input on the final output as described before $r_{i}=\partial f_{\theta}(.)/\partial{\bm{x}}_{i}$ [66].
In context to adversarial examples, Chakraborty et al. [67] observes a consistent shift in the region highlighted in the gradient activation heatmap, reflecting its contribution to decision-making across all models under adversarial attacks.
@@ 3 DEFINITIONS IN OOD DETECTION
Table 2: Overview of ID and OOD definition from several OOD detectors. The post-hoc detectors are divided into the following categories [5]: Classifiction-based, Density-based, Distance-based. We also mark: Supervised and Adversarial Robust, which are not post-hoc detectors.
@@
MDS [11] CIFAR-10 SVHN, TinyImageNet, LSUN, Adversarial Examples DenseNet, ResNet
@@
The robustness definition in the field of OOD detectors has been ambiguous when it comes to attack methods. There are two categories of adversarial examples. The first one merely attacks the underlying pre-trained classifier and the second one aims to fool the OOD detector itself. Adversarial robustness can be considered in a classifier (Unified Robustness) or the OOD detector (Robust OOD Detection) according to [77]. This work focuses on unified robustness, which belongs to the covariate shift. For image classification, a dataset $\mathcal{D}=\{({\bm{x}}_{i},y_{i});{\bm{x}}_{i}\in{\mathcal{X}},y_{i}\in{\mathcal{Y}}\}$ sample from a training distribution $\hat{P}_{\rm{data}}({\bm{x}},y)$ is used to train some classifier $C:{\mathcal{X}}\rightarrow{\mathcal{Y}}$. In real-world deployments, distribution shift occurs when classifier $C$ receives data from test distribution $\hat{P}_{\rm{test}}({\bm{x}},y)$ where $\hat{P}_{\rm{data}}({\bm{x}},y)\neq\hat{P}_{\rm{test}}({\bm{x}},y)$ [78]. An OOD detector is a scoring function $s$ that maps an image ${\bm{x}}$ to a real number $\mathbb{R}$ such that some threshold $\tau$ arrives at the detection rule $f({\bm{x}}):\text{ID if }s({\bm{x}})\geq\tau\text{, OOD otherwise}$.
@@
In Table 2, we present a comparison of the selected ID and OOD datasets utilized by each OOD detector. Our focus lies on experiments employing the most extensive available datasets. Additionally, we enrich this analysis by including adversarial robust OOD detectors discussed in Section 2.3. Notably, recent post-hoc detectors predominantly utilize ImageNet-1K as their standard ID datasets, whereas adversarial robust OOD detectors are more commonly evaluated on the smaller CIFAR-10 dataset. Furthermore, the architecture of adversarial robust methods diverges from lightweight post-hoc designs, often relying on adversarial training and auxiliary data, as elaborated in Section 2.3. The popular adversarial ODD are the FGSM and PGD attacks. Table 2 gives an overview of several detectors as well as considered ID and OOD datasets and model architectures. The ImageNet-1K dataset together with ResNet-50 architecture has become standard for ID. Popular OD datasets are iNaturalist, SUN, Places, and Textures. Some OOD detectors, such as ALOE, OSAD, ADT, and ATOM (see Section 2.3), aim to be robust adversarial. They usually take evading attacks (see Section 2.1) such as FGSM or PGD as OOD. These computationally expensive methods only show empirical results on the small-scale CIFAR-10. A robust OOD detector is built to distinguish whether a perturbed input is OOD. Standardized OOD benchmark frameworks, i.e. OpenOOD [5] or RoboDepth [6] do not currently include unified robustness in their benchmarks. Consequently, both frameworks, give a false sense of encompassing open-world capabilities. They focus on natural distribution [14], where OOD detection in large-scale semantic space has attracted increasing attention [71]. Some OOD datasets have issues, where ID classes are part of the OD dataset [79]. Recently, [80] found a clean semantic shift dataset that minimizes the interference of covariate shift. The experiments show that sota OOD detectors are more sensitive to covariate shifts, and the advances in semantic shift detection are minimal.
@@
Investigating adversarial examples could gain insights into understanding the covariate shift and towards a generalized OOD detection [4]. The difference between benign ${\bm{x}}$ and the attacked counterpart ${\bm{x}}^{\prime}$ is the different attention of the pre-trained classifier $C$ per sample. Let us define an attention map $A$ [81]: The input image is passed through a classifier $C$ to obtain a feature map ${\bm{F}}$. A possible tool could be Grad-CAM [64] (see also Section 2.4) to investigate the attention change between benign ${\bm{x}}$ and adversarial example ${\bm{x}}^{\prime}$ on ${\bm{F}}$ [60] (compare Section 4.2).
@@
So adversarial examples are not from a completely different distribution, but rather adversarial perturbed versions of ID data.

### v5

@@ Abstract
Detecting out-of-distribution (OOD) inputs is critical for safely deploying deep learning models in real-world scenarios. In recent years, many OOD detectors have been developed, and even the benchmarking has been standardized, i.e. OpenOOD. The number of post-hoc detectors is growing fast. They are showing an option to protect a pre-trained classifier against natural distribution shifts and claim to be ready for real-world scenarios. However, its effectiveness in dealing with adversarial examples (AdEx) has been neglected in most studies. In cases where an OOD detector includes AdEx in its experiments, the lack of uniform parameters for AdEx makes it difficult to accurately evaluate the performance of the OOD detector. The existing adversarial OOD definition only declares that AdEx is semantic preserving. We extend this definition by including the attention maps of the underlying neural network classifier and develop a Grad-CAM-based metric to measure the attention shift on the distributional level. This paper investigates the adversarial robustness of 16 post-hoc detectors against various evasion attacks. It also discusses a roadmap for adversarial defense in OOD detectors that would help adversarial robustness. We believe that level 1 (AdEx on a unified dataset) should be added to any OOD detector to see the limitations. The last level in the roadmap (defense against adaptive attacks) we added for integrity from an adversarial machine learning (AML) point of view, which we do not believe is the ultimate goal for OOD detectors.
Code: github.com/adverML/AdvOpenOOD
@@ 1 INTRODUCTION
Adversarial robustness in the context of out-of-distribution (OOD) detection refers to the ability of a detector to correctly identify OOD samples even when they have been adversarial perturbed to evade a deep neural network (DNN). Evasion attacks, e.g. PGD [1], which are designed to fool deep learning classifiers, are difficult to spot as an outlier for OOD detectors. In general, detecting adversarial examples (AdEx) is a challenging task and is almost as hard as classifying them [2]. To prevent errors in real-world applications, it is crucial to detect OOD cases, not only for natural distribution shifts [3] but also for AdEx [4] without degrading the generalizability of the underlying pre-trained classifier [5].
Current standardized benchmarks such as OpenOOD [6] and RoboDepth [7] merely focus on natural distribution shifts and corruptions [8]. Especially OpenOOD aims to make a fair comparison across methods initially developed for anomaly detection [9], model uncertainty [10], open-set recognition (OSR) [11], and OOD detection [12] but neglecting AdEx.
OpenOOD benchmark suite [5] evaluates methods on semantic shift (e.g., samples that are semantically different from the training data, representing truly novel or unseen concepts.) [13] and covariate shift (e.g. samples that come from a different distribution than the training data, but still belong to the same semantic categories) [6].
@@
Current OOD detection methods, as some listed in Table 1, achieve outstanding results on prominent OOD benchmarks, such as the OpenImage-O [14], ImageNet-O [15], Texture [16], and iNaturalist [17, 18]. OOD detection is a very quickly growing field due to the number of methods added to OpenOOD. More specific post-hoc methods with their plug-and-play capabilities on pre-trained classifiers are more flexible and scalable compared to methods that require full retraining on new OOD data [19, 20]. Simple post-hoc methods like KNN [21] are highlighted maintaining good performance on toy-datasets (e.g. MNIST [22], CIFAR-10 or CIFAR-100 [23]), and also show outstanding performance on the more realistic dataset like ImageNet [24] according to [19]. These experiments do neglect the adversarial robustness of “state-of-the-art” (sota) detectors and the real-world capabilities [25] are questionable as past studies had shown [4, 26, 27, 28]. OOD detectors should be able to reliably detect inputs that come from a different distribution than the training data, even if the shift is subtle or complex. AdEx remains challenging because they share the same semantics as the training data, but aim to modify the classifier’s output.
@@
  • We revise the definition of adversarial OOD regarding post-hoc OOD methods to finally have a common understanding of robust adversarial OOD detection referring to Section 3. Finally, we extended the existing definition with attention maps (see Equation 1).
  • We examine 16 post-hoc OOD detectors by delving into their current ability to detect AdEx — an aspect that has been disregarded.
  • We expand the OpenOOD framework with evasive attacks and provide adversarial OOD datasets to have standardized attack parameters. Moreover, we analyze the semantic shift between benign and AdEx by using Grad-CAM combined with a novel metric.
@@ 2 RELATED WORK
This section discusses existing white-box evasion attacks, the benefits of post-hoc detectors, existing methods for detecting AdEx, and finally Grad-CAM, a tool for understanding the neural networks’ attention shift initiated by AdEx.
@@ 2.1 Evasion Attacks Crafting Inliers
The objective of evasion attacks is to generate AdEx that will result in the misclassification of inputs at deep learning models [29]. It is possible to distinguish between two types of attack: black-box attacks [30], where the classifier is queried, and white-box attacks [31], where the network is under the attacker’s complete control. A white-box threat model is strictly stronger. Evasion attacks try to find the smallest possible perturbation, often imperceptible to humans, to manipulate the model’s decision boundaries.
More formally, for an input ${\bm{x}}$ with the ground-truth label $y$, an adversarial example ${\bm{x}}^{\prime}$ is crafted by adding small noise $\bm{\delta}$ to ${\bm{x}}$ such that the predictor model loss ${\bm{J}}({\bm{x}}^{\prime},y)$ is maximized. The $L^{p}$ norm of the adversarial noise should be less than a specified value $\epsilon$, i.e., $\left\|{\bm{x}}-{\bm{x}}^{\prime}\right\|\leq\epsilon$, e.g., $\epsilon=8/255$ [32], to ensure that the image does not change semantically. The attack method Fast Gradient Sign Method (FGSM) by [33] maximizes the loss function in a single step by taking a step towards the sign of the gradient of ${\bm{J}}({\bm{x}},y)$ w.r.t. to ${\bm{x}}$: ${\bm{x}}^{\prime}={\bm{x}}+\epsilon\text{sign}(\nabla_{{\bm{x}}}{\bm{J}}({\bm{x}}^{\prime},y))$, where the noise meets the $L^{\infty}$ norm bound $\epsilon$. Furthermore, this approach can be applied iteratively, as shown by [34], using a reduced step size $\alpha$: $vx^{\prime}_{0}={\bm{x}},\;{\bm{x}}^{\prime}_{t+1}={\bm{x}}^{\prime}_{t}+\alpha\epsilon\text{sign}(\nabla_{{\bm{x}}}{\bm{J}}({\bm{x}}^{\prime}_{t},y))$, where in each step, the perturbation should be constrained within the $L^{\infty}$ ball of radius $\epsilon$. This constraint is a characteristic of the Projected Gradient Descent (PGD) attack proposed by [1], which is commonly considered a standard attack for the evaluation of model robustness [35]. Masked PGD (mPGD) [36] is a variant of the PGD attack that restricts perturbations to a specific region within an image. The PGD attack is extended towards a patch: ${\bm{x}}_{t+1}^{\prime}=\text{Clip}_{{\bm{x}}^{\prime}}\left({\bm{x}}_{t}^{\prime}+\alpha\cdot\text{sign}(\nabla_{{\bm{x}}}{\bm{J}}({\bm{x}}_{t}^{\prime},y,\bm{\theta})[patch]),\epsilon\right).$ In this context, the term “patch” refers to the region specified as $[x:x+h,y:y+w]$, where $[x,y,h,w]$ are the provided patch’s coordinates and dimensions. AutoAttack (AA) [37] consists of four attacks, with the first two being variants of PGD. It remains uncertain whether AA is superior to PGD [38, 39, 40]. Recently, [41] modified PGD, which could also enhance the effectiveness of AA and show that PGD is an essential part of AA.
@@ 2.2 Advantages of Post-Hoc OOD Detectors
Moreover, post-hoc OOD methods can be augmented with other techniques, such as those employed in OSR [46] or uncertainty estimation [47]. Combining different techniques means also the post-hoc methods become more complex and thus the attack surface might become larger, i.e. attacks against uncertainty estimation [10] differ from the evasion attacks. It is not necessarily the case that a post-hoc method must be combined with other techniques. This demonstrates SAFE [48].
@@ 2.3 OOD Adversarial Detection
Ensuring the protection of deployed DL models is the aim of OOD detectors. Merely detecting [2] is already a challenging, but the task of providing comprehensive defense [49] against unknown threats is challenging. Every defense mechanism can be circumvented at some point [50]. Many OOD detectors can be easily evaded by slightly perturbing benign OOD inputs, creating OOD AdEx that reveals a severe limitation of current open-world learning frameworks [51, 52]. Even adversarial training-based defense methods, effective against ID adversarial attacks, struggle against OOD AdEx [52]. In the past years, various defensive techniques [53] and combinations of them have surfaced to combat threats such as adversarial training [1, 54, 55], gradient masking for obfuscation [56], input transformations such as input purification [57, 58]. However, attackers consistently adapt their adversarial attacks to the specific defense mechanisms [59]. According to [60], optimization-based defenses could be a promising future, because they can adapt during test-time towards the input.
@@
2.4 CAM-based Explanations and Consistent Shift by AdEx
Explainable AI (XAI) [63] can be used as a defense mechanism [64, 65, 66] to spot AdEx [67]. The Grad-CAM heatmap provides a clear indication, for a given image, of the regions that the Convolutional Neural Network (CNN) focuses on to make a decision. Class Activation Maps (CAMs) can be considered saliency maps specific to the input [68]. These maps emerge from the accumulated and up-scaled activations at a particular convolutional layer, typically the penultimate layer. The classification is estimated through a linear combination of the activation of units in the final layer(s) of the feature selection network: $f_{\theta}(.)\approx\sum_{i}\sum_{k}w_{k}a_{ki},$ where $a_{ki}$ represents the activation of the k-th channel of unit $i$, and $w_{k}$ denotes the learned weights. The relevance values are subsequently defined as $r_{i}=\sum_{k}w_{k}a_{ki}$. The determination of these weights depends on the specific CAM variant employed [69, 70, 71]. Grad-CAM weights the activations using gradients: $w_{K}:=\frac{\partial f_{\theta}(.)}{\partial a_{ki}}.$ This weighting directly links to more fundamental explanations that merely estimate the influence of the input on the final output as described before $r_{i}=\partial f_{\theta}(.)/\partial{\bm{x}}_{i}$ [72].
In context to AdEx, [73] observes a consistent shift in the region highlighted in the gradient activation heatmap, reflecting its contribution to decision-making across all models under adversarial attacks.
@@ 3 DEFINITIONS IN OOD DETECTION
Table 2: Overview of ID and OOD definition from several OOD detectors. The post-hoc detectors are divided into: Classifiction-based, Density-based, Distance-based [6]. We also mark: Supervised and Adversarial Robust, which are not post-hoc detectors.
@@
MDS [12] CIFAR-10 SVHN, TinyImageNet, LSUN, AdEx DenseNet, ResNet
@@
The robustness definition in the field of OOD detectors has been ambiguous when it comes to attack methods. There are two categories of AdEx. The first one merely attacks the underlying pre-trained classifier and the second one aims to fool the OOD detector itself. Adversarial robustness can be considered in a classifier (Unified Robustness) or the OOD detector (Robust OOD Detection) according to [83]. This work focuses on unified robustness, which belongs to the covariate shift. For image classification, a dataset $\mathcal{D}=\{({\bm{x}}_{i},y_{i});{\bm{x}}_{i}\in{\mathcal{X}},y_{i}\in{\mathcal{Y}}\}$ sample from a training distribution $\hat{P}_{\rm{data}}({\bm{x}},y)$ is used to train some classifier $C:{\mathcal{X}}\rightarrow{\mathcal{Y}}$. In real-world deployments, distribution shift occurs when classifier $C$ receives data from test distribution $\hat{P}_{\rm{test}}({\bm{x}},y)$ where $\hat{P}_{\rm{data}}({\bm{x}},y)\neq\hat{P}_{\rm{test}}({\bm{x}},y)$ [84]. An OOD detector is a scoring function $s$ that maps an image ${\bm{x}}$ to a real number $\mathbb{R}$ such that some threshold $\tau$ arrives at the detection rule $f({\bm{x}}):\text{ID if }s({\bm{x}})\geq\tau\text{, OOD otherwise}$.
@@
We present a comparison of the selected ID and OOD datasets utilized by each OOD detector from Table 1 in Table 2. Our focus lies on experiments employing the most extensive available datasets. Additionally, we enrich this analysis by including adversarial robust OOD detectors discussed in Section 2.3. Notably, recent post-hoc detectors predominantly utilize ImageNet-1K as their standard ID datasets, whereas adversarial robust OOD detectors are more commonly evaluated on the smaller CIFAR-10 dataset. Furthermore, the architecture of adversarial robust methods diverges from lightweight post-hoc designs, often relying on adversarial training and auxiliary data, as elaborated in Section 2.3. The popular adversarial ODD are the FGSM and PGD attacks. The Table 2 gives an overview of several detectors as well as considered ID and OOD datasets and model architectures. The ImageNet-1K dataset together with ResNet-50 architecture has become standard for ID. Popular OD datasets are iNaturalist, SUN, Places, and Textures. Some OOD detectors, such as ALOE, OSAD, ADT, and ATOM (see Section 2.3), aim to be robust adversarial. They usually take evading attacks (see Section 2.1) such as FGSM or PGD as OOD. These computationally expensive methods only show empirical results on the small-scale CIFAR-10. A robust OOD detector is built to distinguish whether a perturbed input is OOD. Standardized OOD benchmark frameworks, i.e. OpenOOD [6] or RoboDepth [7] do not currently include unified robustness in their benchmarks. Consequently, both frameworks, give a false sense of encompassing open-world capabilities. They focus on natural distribution [15], where OOD detection in large-scale semantic space has attracted increasing attention [77]. Some OOD datasets have issues, where ID classes are part of the OD dataset [85]. Recently, [86] found a clean semantic shift dataset that minimizes the interference of covariate shift. The experiments show that sota OOD detectors are more sensitive to covariate shifts, and the advances in semantic shift detection are minimal.
@@
Investigating AdEx could gain insights into understanding the covariate shift and towards a generalized OOD detection [5]. The difference between benign ${\bm{x}}$ and the attacked counterpart ${\bm{x}}^{\prime}$ is the different attention of the pre-trained classifier $C$ per sample. Let us define an attention map $A$ [87]: The input image is passed through a classifier $C$ to obtain a feature map ${\bm{F}}$. A possible tool could be Grad-CAM [70] (see also Section 2.4) to investigate the attention change between benign ${\bm{x}}$ and adversarial example ${\bm{x}}^{\prime}$ on ${\bm{F}}$ [66] (compare Section 4.2).
@@
So AdEx are not from a completely different distribution, but rather adversarial perturbed versions of ID data.